OpenStudy (campbell_st):

find the locus of a point P that moves such that the distance of P from the lines 3x - 4y + 1 = 0 and 12x +5y + 3 = 0 is in the ratio 3:1

OpenStudy (perl):

there is a formula for distance from a point to a line , you might be able to use that

OpenStudy (campbell_st):

oh, you mean this one \[d = \frac{\left| Ax + By + C \right|}{\sqrt{A^2 + B^2}}\]

OpenStudy (perl):

the formula is distance ( xo, yo ) ( ax + by + c = 0 ) = | axo + byo + c | / sqrt(a^2 + b^2 )

OpenStudy (perl):

yes :)

OpenStudy (campbell_st):

where the point is P(x, y)

OpenStudy (perl):

so this might work solve | 3x - 4y + 1 | / sqrt( 3^2 + 4^2 ) = 3 * | 12x + 5y + 3 | / sqrt( 12^2 + 5^2 )

OpenStudy (perl):

the locus is a pair of intersecting lines

OpenStudy (campbell_st):

so its then \[\frac{13(3x - 4y + 1)}{3} = \frac{5(12x + 5y + 3)}{1}\] or \[(39x - 52y + 13) = 3(50x + 25y + 15)\] so then you would say if the point is between the lines then use 3 and collect line terms then is the lines are in the same side of the point you would use -3 since its the external division in a ratio is that what you're saying..?

OpenStudy (campbell_st):

oops should read \[(39x - 52y + 13) = 3(60x + 25y + 15)\]

OpenStudy (perl):

almost, there are two cases because of the absolute value

OpenStudy (perl):

oh i think you considered both cases

OpenStudy (campbell_st):

so what I wrote is what you are saying.... the point is between the lines... and the lines can be on the same side of the point....

OpenStudy (perl):

right

OpenStudy (campbell_st):

wow... thanks for the help..

OpenStudy (perl):

also it would be nice to be able to test this , plug in points

OpenStudy (campbell_st):

maybe you you make a geogebra applet... so show it

OpenStudy (perl):

yeah that would be cool :)

OpenStudy (campbell_st):

I'll ask my teacher if he can do it

OpenStudy (campbell_st):

I read about another side so posted it on there... and people gave an elipseand circle as the answer... I knew that didn't look right

OpenStudy (perl):

do you have a link to that

OpenStudy (campbell_st):

it was study pool... tutors can get paid...

OpenStudy (perl):

its difficult to test because I get fractions when I solve for y

OpenStudy (campbell_st):

why solve for y... its a locus for the values in the point keep changing and are described by the 2 linear equations...

OpenStudy (perl):

i want to plot all three lines and test it

OpenStudy (campbell_st):

all that would acheive is showing a specific P

OpenStudy (perl):

right

OpenStudy (perl):

and you can move P around

OpenStudy (campbell_st):

I'd need to look at it and the point P in the solution line can be moved

OpenStudy (campbell_st):

anyway.... that's my Year 11 homework done... thanks

OpenStudy (perl):

:)

OpenStudy (perl):

I just tested it for the point P ( 1, -277/23) , and it works. I used the negative line http://www.wolframalpha.com/input/?i=simplify+-%28+3x+-+4y+%2B+1+%29+%2F5+%3D+3+%28+12x+%2B+5y+%2B+3+%29+%2F13

OpenStudy (perl):

distance [( 1, -277/23) , ( 3x -4y + 1 = 0 ) ] = 240/23 distance [(1,-277/23) , ( 12x + 5y + 3 = 0 ) ] = 80/23